How much time is needed to measure kinetic energy of an electron whose speed is 10 m/s with an uncertainty of not more than 0.1% How much the electron will travel in this period of time?
1
Expert's answer
2019-08-14T09:47:57-0400

1) From the uncertainty principle:


"∆E∆t ≥ \frac{\hslash}{2}"

"\frac{\Delta E}{E}=\frac{\hslash}{2E\Delta t}=0.001"

"\Delta t=\frac{\hslash}{0.001m_e v^2}=1.16\ ms"

2)


"x=v\Delta t=(10)(0.00106)=0.0106\ m=1.06\ cm"